# Floorplan-SLAM: A Real-Time, High-Accuracy, and Long-Term Multi-Session Point-Plane SLAM for Efficient Floorplan Reconstruction

### v2

@@ Abstract
[…]
@@ III-B Plane Landmark Management
For any two plane landmarks $\mathbf{\Pi}_{i}$ and $\mathbf{\Pi}_{j}$, if the angle between their normal vectors is less than $\theta_{c}$ and the minimum point-to-plane distances from the support points of $\mathbf{\Pi}_{i}$ to $\mathbf{\Pi}_{j}$ and from $\mathbf{\Pi}_{j}$ to $\mathbf{\Pi}_{i}$ are both less than $d_{c}$, we remove the plane landmark with fewer keyframe observations and transfer its observation information and support points to another plane landmark, followed by RANSAC to update the plane points. The above operations continue throughout the entire process of the system. In our experiments, $d_{a}$ is set to 1 cm, $\theta_{c}$ is set to 5, and $d_{c}$ is set to 2 cm.
@@ III-C Multi-Session Map Merging
We use DBoW2 [17] for place recognition and compute the aligning transformation $\mathbf{T}_{cm}$ between the two maps. Then, a local window is defined, comprising matching keyframes from both maps, their respective neighbors in the covisibility graph, and the observed point and plane landmarks. Duplicated landmarks from the current map are removed within this local window, and their observations are transferred to the corresponding landmarks in the matching map. A local bundle adjustment (BA) is performed to optimize all keyframes and the point and plane landmarks they observe within the local window. Pose graph optimization is then conducted on the essential graph of the merged map, with the keyframes in the local window fixed. Finally, a global BA is executed in a separate thread to further optimize the merged map and enhance overall consistency. By merging sequences captured at different times, we do not need to capture all map areas in a single pass, which is crucial for large-scale scene long-term mapping, as it significantly reduces the computational and storage burden on the acquisition devices. Notably, during each BA process, the parameters of plane landmark $\mathbf{\Pi}_{k}$ are optimized, and its support points are updated by aggregating the support points of its observations $\boldsymbol{\pi}_{i,k}$ in the keyframes into $\mathbf{\Pi}_{k}$ followed by RANSAC, thereby providing a more accurate scene layout for the subsequent floorplan reconstruction.
@@ V-B Localization Accuracy Evaluation
Accurate poses allow planes extracted from different frames to be seamlessly integrated into the global map, maintaining the global consistency of the incrementally reconstructed floorplan. We evaluate the localization performance of our method in large-scale indoor scenes and compare it with the most relevant state-of-the-art visual SLAM systems, including ORB-SLAM3 (stereo mode) [19], Stereo-Plane-SLAM [11], and RSS [10]. To assess the impact of plane features, we also conduct an ablation study using a point-only variant of our system. Experiments are performed on the public VECtor dataset [20] and a self-collected dataset, both containing sequences exceeding 100 meters. The self-collected dataset was captured using a ZED2 camera with GeoSLAM ZEB as ground truth (Fig. 5). We align the estimated trajectory with ground truth using the Umeyama algorithm without scaling and use the RMSE of the absolute trajectory error (ATE) to measure global drift. Each sequence is run ten times to mitigate multi-thread randomness, and the median results are reported.
@@
Corridor 1.53 1.57 1.38 1.54 1.29
Room 1.40 1.25 1.18 1.38 1.07
Stair 2.35 1.79 1.66 2.40 1.60
@@ V-C Floorplan Reconstruction Performance Evaluation
We evaluate the floorplan reconstruction performance of our method on the VECtor [1] and a self-collected dataset, comparing it with two state-of-the-art offline floorplan reconstruction methods: VecIM [1] and RoomFormer [6]. The former employs a gravity-aligned point cloud as input, while the latter uses a 2D density map obtained by projecting the point cloud along the gravity axis. In contrast, our method uses only stereo images as input. VecIM and RoomFormer both utilize a dense multi-view stereo (MVS) point cloud reconstructed by COLMAP [21], which takes advantage of more comprehensive global optimization and additional multi-view information. This results in a more accurate point cloud than that generated by our SLAM system, ensuring fairness in the comparison. We uniformly sample 10K points from the reconstructed floorplan and compute the Hausdorff distance between these sampled points and the ground-truth point cloud to evaluate the reconstruction accuracy, using a dense, high-precision LiDAR point cloud as the ground truth.
@@ V-C1 Evaluation on the VECtor Dataset
Figure 7: Floorplan reconstruction results on the self-collected dataset of VecIM, RoomFormer and Ours. For each reconstruction result, the mean Hausdorff error (m) and the error model are computed from the output to the ground truth. From top to bottom, we display the office, café, and meeting rooms, respectively. Our method achieves the best performance in both reconstruction accuracy and regularity.
@@ V-C2 Evaluation on the Self-collected Dataset
To evaluate the floorplan reconstruction performance of our method in complex indoor environments containing numerous non-lifelong structures (e.g., tables, chairs), and to assess its long-term floorplan reconstruction capability, we used a ZED2 camera to collect data in three large-scale scenes, covering meeting rooms, offices, and a café, among other places, each covering an area of 1000 square meters. In the meeting rooms scenario, due to the inconsistent accessibility of different conference rooms, we conducted five separate data collections over two days, with each sub-session covering a different region of the scene, as shown in Fig. 1. The qualitative and quantitative results of floorplan reconstruction on the self-collected dataset are presented in Fig. 7. Our method still reconstructs the most regular and faithful floorplan. Benefiting from rich textures, the MVS point cloud exhibits relatively low error, allowing VecIM to accurately reconstruct main structures. However, VecIM is unable to focus solely on extracting lifelong structures and detects numerous temporary structures such as table sides, adding redundant elements and failing to handle navigable areas effectively. In contrast, although our method may initially detect planes from non-lifelong objects, these planes typically do not persist across multiple frames and are not considered valid plane landmarks for the floorplan reconstruction module, enabling a focus on lifelong structures. Moreover, by fully leveraging trajectory information, we effectively handle openings in navigable areas, such as corridors and room entrances, ensuring that the reconstructed floorplan accurately represents passable regions. In addition, although our self-collected dataset contains room scenes, RoomFormer lacks generalizability for large-scale environments and thus can only generate a limited number of correctly identified rooms. Notably, by sequentially merging multiple subsequences, our method yields a complete large-scale floorplan that encompasses multiple meeting rooms, exhibiting remarkable accuracy and regularity, which strongly demonstrates the long-term globally consistent mapping capability of our approach. As a result, there is no need to acquire the complete scene data in a single pass, greatly reducing the complexity of data collection and reconstruction.
@@ V-D Runtime Analysis
The runtime analysis was conducted on a desktop with an AMD Ryzen 7 4800H CPU (running at 2.9 GHz) and 32 GB of RAM. We evaluated our method on the school sequence from the VECtor dataset and on the meeting rooms sequence from our self-collected dataset, comparing the results with ORB-SLAM3. Table III shows the runtimes of the two threads where our method and ORB-SLAM3 differ most; the other threads show comparable execution times. We extract Sobel and ORB features in parallel CPU threads, with the former being faster. Moreover, ORB extraction is the main time-consuming step compared to other plane extraction operations, such as mesh generation, clustering, and plane parameter estimation. As a result, our overall plane extraction time is close to that of ORB feature extraction. The floorplan reconstruction module comprises two primary components: segment generation and selection. We evaluate the runtime by measuring the duration of the final reconstruction in each sequence, which is the most time-consuming due to the largest number of segments. Notably, considering that valid plane landmarks are typically updated every several frames, we perform floorplan reconstruction every five frames. This frequency is sufficient to provide timely updates, thereby enabling real-time floorplan reconstruction. We also recorded the total reconstruction time consumed by VecIM, including the MVS dense point cloud reconstruction and the floorplan reconstruction stages. The first stage took 16 hours and 38.9 minutes, and the second stage took 5.1 minutes, resulting in a total of 16 hours and 44 minutes. In contrast, we processed only stereo images at 45 FPS, and completed the entire floorplan reconstruction in 9.4 minutes, which is two orders of magnitude faster than VecIM. This result strongly demonstrates the efficiency and practicality of our method.
@@ VI Conclusion
[…]

### v3

@@ Abstract
Floorplan reconstruction provides structural priors essential for reliable indoor robot navigation and high-level scene understanding. However, existing approaches either require time-consuming offline processing with a complete map, or rely on expensive sensors and substantial computational resources. To address the problems, we propose Floorplan-SLAM, which incorporates floorplan reconstruction tightly into a multi-session SLAM system by seamlessly interacting with plane extraction, pose estimation, back-end optimization, and loop & map merging, achieving real-time, high-accuracy, and long-term floorplan reconstruction using only a stereo camera. Specifically, we present a robust plane extraction algorithm that operates in a compact plane parameter space and leverages spatially complementary features to accurately detect planar structures, even in weakly textured scenes. Furthermore, we propose a floorplan reconstruction module tightly coupled with the SLAM system, which uses continuously optimized plane landmarks and poses to formulate and solve a novel optimization problem, thereby enabling real-time and high-accuracy floorplan reconstruction. Note that by leveraging the map merging capability of multi-session SLAM, our method supports long-term floorplan reconstruction across multiple sessions without redundant data collection. Experiments on the VECtor and the self-collected datasets indicate that Floorplan-SLAM significantly outperforms state-of-the-art methods in terms of plane extraction robustness, pose estimation accuracy, and floorplan reconstruction fidelity and speed, achieving real-time performance at 25–45 FPS without GPU acceleration, which reduces the floorplan reconstruction time for a 1000 $\mathrm{m}^{2}$ scene from 16 hours and 44 minutes to just 9.4 minutes.
@@ III-B Plane Landmark Management
We create new plane landmarks using planes extracted from keyframes that do not match any existing plane landmarks in the map, where planes are considered matched if the angle between their normal vectors is less than $\theta_{m}$ and the minimum distance from a planar point to the plane is less than $d_{m}$. These newly created plane landmarks are initially marked as invalid due to their potential unreliability, and become valid once they have been observed by more than thirty frames and three keyframes. Only valid plane landmarks are utilized for pose estimation and floorplan reconstruction. However, if a valid plane landmark is observed by at most one keyframe due to excessive optimization errors, it is reverted to invalid. For any two plane landmarks $\mathbf{\Pi}_{i}$ and $\mathbf{\Pi}_{j}$, if the angle between their normal vectors is less than $\theta_{m}$ and the minimum point-to-plane distances from the support points of $\mathbf{\Pi}_{i}$ to $\mathbf{\Pi}_{j}$ and from $\mathbf{\Pi}_{j}$ to $\mathbf{\Pi}_{i}$ are both less than $d_{m}$, we remove the plane landmark with fewer keyframe observations and transfer its observation information and support points to another plane landmark, followed by RANSAC to update the plane points. The above operations continue throughout the entire process of the system. In our experiments, $\theta_{m}$ is set to 5, and $d_{m}$ is set to 2 cm.
@@ III-C Multi-Session Map Merging
We use DBoW2 [17] for place recognition and compute the aligning transformation $\mathbf{T}_{cm}$ between the two maps. The points and plane landmarks from the current map are transformed into the matching map using $\mathbf{T}_{cm}$. Duplicated landmarks from the current map are removed, and their observations are transferred to the corresponding landmarks in the matching map. We then perform local bundle adjustment (BA), pose graph optimization, and global BA in sequence, similar to [18], to optimize the merged map and enhance overall consistency. By merging sequences captured at different times, we do not need to capture all map areas in a single pass, which is crucial for large-scale scene long-term mapping, as it significantly reduces the computational and storage burden on the acquisition devices. Notably, during each BA process, the parameters of plane landmark $\mathbf{\Pi}_{k}$ are optimized, and its plane points are updated by aggregating the support points of its observations $\boldsymbol{\pi}_{i,k}$ in the keyframes into $\mathbf{\Pi}_{k}$ followed by RANSAC, thereby providing a more accurate scene layout for the subsequent floorplan reconstruction.
@@ V Experiment
We evaluated our method on the public VECtor [20] dataset and a self-collected dataset. The self-collected dataset was captured using a ZED2 camera with a GeoSLAM ZEB as ground truth, as shown in Fig. 5. All experiments were carried out on a desktop with an AMD Ryzen 7 4800H CPU (running at 2.9 GHz) and 32 GB of RAM.
@@ V-B Localization Accuracy Evaluation
Accurate poses allow planes extracted from different frames to be seamlessly integrated into the global map, maintaining the global consistency of the incrementally reconstructed floorplan. We evaluate the localization performance of our method in large-scale indoor scenes and compare it with the most relevant state-of-the-art visual SLAM systems, including ORB-SLAM3 (stereo mode) [18], Stereo-Plane-SLAM [11], and RSS [10]. To assess the impact of plane features, we also conduct an ablation study using a point-only variant of our system. Experiments are performed on the public VECtor dataset [20] and the self-collected dataset, both containing sequences several hundred meters in length. We disable the loop closure module and align the estimated trajectory with ground truth using the Umeyama algorithm without scaling. The RMSE of the absolute trajectory error (ATE) is used to measure global drift. Each sequence is run ten times to mitigate multi-thread randomness, and the median results are reported.
@@
café 1.53 1.57 1.38 1.54 1.29
offices 1.40 1.25 1.18 1.38 1.07
meeting rooms 2.35 1.79 1.66 2.40 1.60
@@ V-C Floorplan Reconstruction Performance Evaluation
We evaluate the floorplan reconstruction performance of our method on the VECtor [1] and the self-collected dataset, comparing it with two state-of-the-art offline floorplan reconstruction methods: VecIM [1] and RoomFormer [6]. The former employs a gravity-aligned point cloud as input, while the latter uses a 2D density map obtained by projecting the point cloud along the gravity axis. In contrast, our method uses only stereo images as input. VecIM and RoomFormer both utilize a dense multi-view stereo (MVS) point cloud reconstructed by COLMAP [21], which takes advantage of more comprehensive global optimization and additional multi-view information. This results in a more accurate point cloud than that generated by our SLAM system, ensuring fairness in the comparison. We uniformly sample 10K points from the reconstructed floorplan and compute the Hausdorff distance between these sampled points and the ground-truth point cloud to evaluate the reconstruction accuracy, using a dense, high-precision LiDAR point cloud as the ground truth.
@@ V-C1 Evaluation on the VECtor Dataset
Figure 7: Floorplan reconstruction results on the self-collected dataset of VecIM, RoomFormer and Ours. For each reconstruction result, the mean Hausdorff error (m) and the error model are computed from the output to the ground truth. From top to bottom, we display the offices, café, and meeting rooms, respectively. Our method achieves the best performance in both reconstruction accuracy and regularity.
@@ V-C2 Evaluation on the Self-collected Dataset
We evaluate the floorplan reconstruction and long-term reconstruction performance of our method on the meeting rooms, offices, and café sequences from our self-collected dataset, all of which are complex indoor environments with numerous non-lifelong structures (e.g., tables, chairs), each spanning an area of approximately 1000 square meters. In the meeting rooms scenario, due to the inconsistent accessibility of different conference rooms, we conducted five separate data collections over two days, with each sub-session covering a different region of the scene, as shown in Fig. 1. The qualitative and quantitative results of floorplan reconstruction on the self-collected dataset are presented in Fig. 7. Our method still reconstructs the most regular and faithful floorplan. Benefiting from rich textures, the MVS point cloud exhibits relatively low error, allowing VecIM to accurately reconstruct main structures. However, VecIM is unable to focus solely on extracting lifelong structures and detects numerous temporary structures such as table sides, adding redundant elements and failing to handle navigable areas effectively. In contrast, although our method may initially detect planes from non-lifelong objects, these planes typically do not persist across multiple frames and are not considered valid plane landmarks for the floorplan reconstruction module, enabling a focus on lifelong structures. Moreover, by fully leveraging trajectory information, we effectively handle openings in navigable areas, such as corridors and room entrances, ensuring that the reconstructed floorplan accurately represents passable regions. In addition, although our self-collected dataset contains room scenes, RoomFormer lacks generalizability for large-scale environments and thus can only generate a limited number of correctly identified rooms. Notably, by sequentially merging multiple subsequences, our method yields a complete large-scale floorplan that encompasses multiple meeting rooms, exhibiting remarkable accuracy and regularity, which strongly demonstrates the long-term globally consistent mapping capability of our approach. As a result, there is no need to acquire the complete scene data in a single pass, greatly reducing the complexity of data collection and reconstruction.
@@ V-D Runtime Analysis
We evaluated our method on the school sequence from the VECtor dataset and on the meeting rooms sequence from our self-collected dataset, comparing the results with ORB-SLAM3. Table III shows the runtimes of the two threads where our method and ORB-SLAM3 differ most; the other threads show comparable execution times. We extract Sobel and ORB features in parallel CPU threads, with the former being faster. Moreover, ORB extraction is the main time-consuming step compared to other plane extraction operations, such as mesh generation, clustering, and plane parameter estimation. As a result, our overall plane extraction time is close to that of ORB feature extraction. The floorplan reconstruction module comprises two primary components: segment generation and selection. We evaluate the runtime by measuring the duration of the final reconstruction in each sequence, which is the most time-consuming due to the largest number of segments. Notably, considering that valid plane landmarks are typically updated every several frames, we perform floorplan reconstruction every five frames. This frequency is sufficient to provide timely updates, thereby enabling real-time floorplan reconstruction. We also recorded the total reconstruction time consumed by VecIM, including the MVS dense point cloud reconstruction and the floorplan reconstruction stages. The first stage took 16 hours and 38.9 minutes, and the second stage took 5.1 minutes, resulting in a total of 16 hours and 44 minutes. In contrast, we processed only stereo images at 45 FPS, and completed the entire floorplan reconstruction in 9.4 minutes, which is two orders of magnitude faster than VecIM. This result strongly demonstrates the efficiency and practicality of our method.
@@ VI Conclusion
This paper presents Floorplan-SLAM, a novel approach for real-time, high-accuracy, and long-term floorplan reconstruction, significantly reducing reconstruction time while maintaining superior accuracy and robustness. The proposed method consistently outperformed state-of-the-art approaches in three key aspects: plane extraction robustness, pose estimation accuracy, and floorplan reconstruction fidelity and speed. Experimental results on the VECtor dataset and our self-collected dataset demonstrate the effectiveness of Floorplan-SLAM. Specifically, our algorithm is able to reconstruct a 1000 $\mathrm{m}^{2}$ floorplan in just 9.4 minutes, compared with the over 16 hours and 44 minutes required by the existing offline method [1], achieving a speedup of two orders of magnitude, while yielding more accurate and structured reconstruction results. The real-time performance, achieving 25–45 FPS without GPU acceleration, further underscores the efficiency and practicality of our approach for large-scale indoor floorplan reconstruction. Future work will focus on deeply integrating the floorplan with the SLAM system to fully leverage the high-level scene understanding it provides, thereby enabling hierarchical localization and map merging in challenging scenarios.